OpenStudy (graciek):

In this unit, you learned about fractions that can be written as terminating decimals and fractions that can be written as nonterminating decimals. Create an addition word problem and a subtraction word problem using fractions. Show how to find the solution to each problem. Once you find the solutions to your problems, rewrite the solutions as either a terminating or nonterminating decimal.

OpenStudy (luckycoins888):

So you write a simple fraction addition problem: 1/2 + 1/5 and a subtraction problem: 1/2 - 1/5 Then you solve them, using what you learned this unit: 1/2+1/5=5/10+2/10=7/10 and 1/2-1/5=5/10-2/10=3/10 And then you convert them into decimals: 7/10 = 7 divided by 10 = 0.7, and 3/10 = 0.3 To get a nonterminating decimal, you use fractions that will result in a fraction such as 22/7, who's answer does not terminate

OpenStudy (luckycoins888):

(medal for fast typing pls XD)

OpenStudy (graciek):

wait so could i put that down for what it says?

OpenStudy (luckycoins888):

Yeh.

OpenStudy (graciek):

ok. thank you :)